OpenStudy (anonymous):

Suppose that a family wants to have four children. (Enter the probabilities as fractions.) (a) What is the sample space? (Let B represent a boy and G represent a girl. To indicate that the first child is a boy, the second child is a boy, the third child is a boy, and the fourth child is a girl enter BBBG. Enter your answers as a comma-separated list.)

OpenStudy (anonymous):

What is the probability of 4 boys? and 4 girls?

OpenStudy (triciaal):

P(4B) =1/2*1/2*1/2*1/2 = (1/2)^4 = P(4G)= same to do the options for 4 children